1 mole of H 2 SO 4 can be written in how many different ways?

1 mole of H2SO4 can be written in how many different ways?

Grade:11

1 Answers

Adarsh poddar
7 years ago
10 ways to represent 1 Mole of H2SO4- 1 Mole of H2SO4- 6.022*10^23 atoms of sulphur-4*6.022*10^23 atoms of oxygen-2*6.022*10^23 atoms of hydrogen- 2+16+32= 50 moles of electrons- 2+16+32= 50 moles of protons- 2+16+32= 50 moles of neutrons-Molar mass = 98 gram/Mole-22.4 litres of H2SO4 -150 moles of protons neutrons and electrons all together.